OpenStudy (help108):

Hello I need help on this one simple question and explain the steps carefully and ill give you medals!!! galore and fans!!!!! 2x - 3 = -6

OpenStudy (instagrammodel):

I'm Here

OpenStudy (help108):

ok its 2x - 3 = -6

OpenStudy (help108):

@Photon335

OpenStudy (wentz):

2x - 3 = -6 add 3 to both sides. now you have 2x = -3 divide by 2 x=-1.5

OpenStudy (help108):

what the freak? u add 3 on both side so 3 plus 2 and then 3 plus 6?

OpenStudy (help108):

when i added 3 to 2x it made 2x? and when i added 3 to -6 it made -3 that makes sense

OpenStudy (help108):

the second part not the first

OpenStudy (wentz):

well since 3 and 6 are both negatives, adding three with cancel out -3, becoming 0 then adding to -6, will be -3. then there are 2 x's on one side, -3 on the other. divide the 2 x's and that -3 by 2 and you have x=-1.5 does that make anymore sense? i feel like i just repeated it in a different way

OpenStudy (help108):

how do u divide the -3 with a 2?

OpenStudy (help108):

so now that it equals -1.5 so then the equation will now be 21.5 - 3 = -6?

OpenStudy (wentz):

nope, it ends at x = 1.5

OpenStudy (help108):

how can this apply to real life? its not a math question i just want to know for myself lol.

OpenStudy (help108):

I understand everything now your awesome dude.! :)

OpenStudy (wentz):

hmmm im not rlly sure bc im really bad at math but this is the only kind im good at and i saw this question and was like pRAISE I CAN HELP

OpenStudy (help108):

okay thanks lol bye!!